OpenStudy (anonymous):

We have : \[\int_0^1 \int_{-\sqrt{4-4y^2}}^{\sqrt{4-4y^2}} \int_{x^2+3y^2}^{4-y^2}dzdxdy\] Task: change the order of integral to dxdzdy

hartnn (hartnn):

since we are only interested in changing the order of x and z, we may totally ignore dy, and the 3rd integral 0 to 1, and treat the 'y' as constant. so, basically you're dealing with, \(\large \int \limits_{-b}^{b} \quad \int \limits_{x^2+a}^{c}dzdx\) does that simlify things ?

OpenStudy (anonymous):

Hmm, it still looks complicated :(

hartnn (hartnn):

ofcourse, after we are done, we will resubstitute for y

hartnn (hartnn):

z=x^2+a is parabola z=c x=-b x=b are 3 lines

hartnn (hartnn):

in x-z plane

OpenStudy (anonymous):